OpenStudy (anonymous):

A regular 20-gon with sides of length 4 units can be divided into 20 equilateral triangles with sides of length 4 units. The area of the 20-gon is ______. (Give the exact form of the answer.)

OpenStudy (anonymous):

well calculate the area of one of the triangles then multiply that area by 20

OpenStudy (anonymous):

you could just use the formula Area of a twentygon = (Base*height/2)*20

OpenStudy (anonymous):

Area of a twentygon = (Base*height/2)*20=160

OpenStudy (anonymous):

4*4/2 *(20)= 16/2*(20)= 8*(20)=160